OpenStudy (anonymous):

Solve: 2x-y=3 4x-3y=7

OpenStudy (anonymous):

i can't help you solve 2x-y=3

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

y=3+2x 4x-3(3+2x)=7 x=1 y=-1

OpenStudy (anonymous):

Multiply 2x-y=3 by -2, so it becomes -4x+2y=-6. -4x+2y=-6 + 4x-3y = 7 = 0 - y = 1 so y= -1. Plug this back in. 2x+1=3 2x=2 x=1.

OpenStudy (anonymous):

y=2x-3 4x-3(2x-3)=7